OpenStudy (liaannne):

You bought a coat at $78.50 and two shirts at $21.99 each. If the sales tax rate is 7%, how much is the total purchase?

563blackghost (563blackghost):

You would need to multiply the cost of each shirt by 2. \(\huge\bf{21.99 \times 2 = 43.98}\) Next we add the coat and shirt costs together. \(\huge\bf{43.98 + 78.50 = 122.48}\) We then would multiply the total by the sales tax `( we must first turn to decimal so it would be 0.07)`. \(\huge\bf{122.48 \times 0.07 = 8.54}\) We would then add it to the total. \(\huge\bf{122.48 + 8.54 =?}\)